OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. y = 4x - 11 y = 4x + 11 y = 4x + 7 y = 4x - 7

OpenStudy (anonymous):

is this flvs? i think i had this question

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what lesson is this ?

OpenStudy (anonymous):

its lesson 3.00. i really don't understand this chapter well.

OpenStudy (anonymous):

k im looking to see if i had this question but are you doing the summer program ?

OpenStudy (anonymous):

is this the pretest?

OpenStudy (scorcher219396):

A formula to find an equation: \[y-y _{a}=m(x-x _{a})\] where ya and xa are the points that you're given, and m is the slope. Then, just plug in your info and simplify y-(-1)=4(x-(-3)) y+1=4(x+3) y+1=4x+12 y=4x+11

OpenStudy (anonymous):

Yes it is

OpenStudy (anonymous):

you know thats only 1 point ? like it wont affect your grade at all

OpenStudy (anonymous):

thank you so much once again